# Analytic Personalized Federated Meta-Learning

### v1

@@ Abstract
Analytic federated learning (AFL) which updates model weights only once by using closed-form least-square (LS) solutions can reduce abundant training time in gradient-free federated learning (FL). The current AFL framework cannot support deep neural network (DNN) training, which hinders its implementation on complex machine learning tasks. Meanwhile, it overlooks the heterogeneous data distribution problem that restricts the single global model from performing well on each client’s task. To overcome the first challenge, we propose an AFL framework, namely FedACnnL, in which we resort to a novel local analytic learning method (ACnnL) and model the training of each layer as a distributed LS problem. For the second challenge, we propose an analytic personalized federated meta-learning framework, namely pFedACnnL, which is inherited from FedACnnL. In pFedACnnL, clients with similar data distribution share a common robust global model for fast adapting it to local tasks in an analytic manner. FedACnnL is theoretically proven to require significantly shorter training time than the conventional zeroth-order (i.e. gradient-free) FL frameworks on DNN training while the reduction ratio is $98\%$ in the experiment. Meanwhile, pFedACnnL achieves state-of-the-art (SOTA) model performance in most cases of convex and non-convex settings, compared with the previous SOTA frameworks.
Index Terms:
gradient-free, analytic federated learning, meta-learning.
I Introduction
@@
Typical FL training schemes include first- [2, 3] (e.g., SGD) and second-order methods [4], which necessitate the computation of gradients and Hessian matrices. Despite the efficiency of these gradient descent (GD)-based methods, their reliance on gradient computations can pose challenges in practice. Specifically, (i) edge devices with poor computational capabilities are incapable of performing frequent gradient calculations. For instance, mobile phones or IoT sensors are primarily designed for inference-only purposes rather than backpropagation [5]. In such cases, only small, pre-trained or pruned models are computationally affordable for execution on these devices [1, 6]. (ii) Gradient information is not always available to support gradient computations. For instance, in situations such as federated hyperparameter tuning [7] and federated black-box attack [8], the gradients cannot be derived due to the non-differentiable nature of the loss function. As a result, GD-based FL frameworks may become infeasible in these scenarios, prompting a strong demand from both academia and industry for the development of gradient-free FL training approaches.
@@
In AFL, the model weights are updated only once and are determined by using closed-form least-square (LS) solutions. Therefore, AFL can achieve a significantly shorter training time than the conventional zeroth-order FL frameworks. However, AFL overlooks the performance limitations of training a single global model for clients with heterogeneous data distributions. In other words, training a unified global model to satisfy the performance requirements of all clients is challenging due to the distinct preferences in data generation or collection by different clients. Personalized FL (PFL) [12] is a promising paradigm to alleviate the heterogeneous data distribution problem. In PFL, each client utilizes the common knowledge, which is the aggregated global model on the PS, to generate a personalized model tailored for its local data distribution. Meta-learning [13, 14, 15], as a representative model optimization technique in PFL, enables the aggregated global model (i.e. the meta-model) to fast adapt to heterogeneous tasks by performing several gradient descents on the global model. This motivates us to introduce meta-learning to AFL to address the heterogeneous data distribution problem in AFL. This introduction results in two challenges: First of all, the current AFL framework [11] does not have the capability of training deep neural networks (DNN), which is the basis of tackling complex machine learning tasks. Thus, this nature makes the introduction of meta-learning to the current AFL framework meaningless when we confront complex machine learning tasks. Secondly, the existing meta-learning method is only suitable for clients using GD-based optimization methods. Therefore, combining meta-learning with AFL directly is impractical since AFL does not involve any gradient approximation or computations.
@@
In this paper, we first resort to ACnnL (analytic convolutional neural network learning) [16], an analytic single-device learning method for DNN, and propose its federated implementation, called FedACnnL, to enable DNN training for AFL. Specifically, for the PS, a layer-wise global model update rule using closed-form LS solutions is designed with a specialized aggregation strategy to realize the layer-wise collaborative DNN training. For the clients, an adaptive batch size algorithm is incorporated to combat the straggler problem of FL, in which the PS must keep idle until all clients finish their local computations. By increasing the local batch size to reduce the local computation time of lagging clients, the idle time of the PS is diminished. Thanks to the weight-invariant property of FedACnnL, this incorporation achieves acceleration without any model performance degradation. Therefore, FedACnnL, as an AFL framework, realizes computation-efficient collaborative DNN training. Then, based on FedACnnL, we propose an analytic personalized federated meta-learning (APFML) framework, called pFedACnnL. A gradient-free meta-learning method is designed in it to combat the heterogeneous data distribution problem. Specifically, the PS groups the clients with similar data distribution to generate a meta-model that contains common knowledge within each group by following the standard training process of FedACnnL. Next, each client can achieve its personalized model tailored for its local data distribution without using any gradient information by analytically solving a local PFL objective. This objective bridges the gap between the personal knowledge of its local dataset and the common knowledge of its meta-model. Our contributions can be summarized as follows:
@@
  • We first resort to ACnnL and propose FedACnnL to enable computation-efficient DNN training in AFL. With a competitive model performance on DNN, FedACnnL is proven to require a greatly shorter training time than the conventional zeroth-order (i.e. gradient-free) FL frameworks both theoretically and empirically.
@@
  • We validate the efficiency of our frameworks on a CPU-cluster-based testbed, which can emulate the environmental settings of data heterogeneity and resource heterogeneity in FL simultaneously. Experimentally, FedACnnL can reduce the training time by $88\%\sim 98\%$ compared with the conventional first- and zeroth-order FL frameworks. Meanwhile, pFedACnnL achieves state-of-the-art (SOTA) model performance in most cases of convex and non-convex settings, compared with other previous SOTA first- and zeroth-order PFL frameworks.
TABLE I: Table of partial notations.
@@
II Preliminaries
II-A Notations
[…]
II-B Analytic Convolutional Neural Network Learning (ACnnL)
@@
Given an $N$-sample dataset $\mathcal{S}=\{\textbf{x}_{n},\textbf{y}_{n}\}_{n=1}^{N}$ where $\textbf{x}_{n}\in{\mathbb{R}}^{d_{0}}$ and $\textbf{y}_{n}\in{\mathbb{R}}^{d_{y}}$ denote the data vector and the label vector of the $i$-th sample respectively. We stack these samples to generate the data matrix $\textbf{X}=\left[\textbf{x}_{1};\textbf{x}_{2};\cdots;\textbf{x}_{N}\right]\in\mathbb{R}^{N\times{d_{0}}}$ and the label matrix $\textbf{Y}=\left[\textbf{y}_{1};\textbf{y}_{2};\cdots;\textbf{y}_{N}\right]\in\mathbb{R}^{N\times{d_{y}}}$. In the $l$-th iteration, the pseudo label matrix of the $l$-th hidden layer $\bar{\textbf{Z}}_{l}$ is generated as $\bar{\textbf{Z}}_{l}=\textbf{YQ}_{l}$, where ${\textbf{Q}_{l}}\in{\mathbb{R}^{d_{y}\times{d_{l}}}}$ is the label encoding matrix which is randomly initialized from a normal distribution. The encoding is linear so that $\bar{\textbf{Z}}_{l}$ can fit the dimensionality of $\textbf{Z}_{l}\in\mathbb{R}^{N\times{d_{l}}}$, which is the real output matrix before entering the activation function of the $l$-th hidden layer $f_{l}(\cdot)$ when X is the input of the DNN model being trained. Specially, $\bar{\textbf{Z}}_{l}=\textbf{Y}$ when training the output layer of the model. Then, ACnnL formulates the supervised learning of the $l$-th layer as an least-square (LS) problem with $l_{2}$ regularization:
@@
where $\textbf{X}_{l}\in\mathbb{R}^{N\times{d_{l-1}}}$ and $\textbf{W}_{l}\in\mathbb{R}^{d_{l-1}\times{d_{l}}}$ denote the real input matrix and model weights of the $l$-th hidden layer respectively. Specially, $_{l}=\textbf{X}$ X when training the input layer. $\gamma$ is the regularization coefficient which is a hyperparameter to control the search space of the LS solutions. Next, we can take the closed-from LS solutions of the weights in $l$-th hidden layer by putting the derivatives with respect to $\textbf{W}_{l}$ on the objective function in the equation 1 to 0. We present the solutions at a sample level as follows:
@@ III Design of FedACnnL
III-A Distributed LS Problem
Based on the supervised learning property of ACnnL in each hidden layer, we can model the weight update problem in each layer as a distributed LS problem. Specifically, suppose we have the aforementioned $N$-sample dataset $\mathcal{S}$ in the section II-B, and $C$ clients of edge devices. The dataset is disjointly distributed onto the $C$ clients, i.e. $\mathcal{S}=\cup_{i=1}^{C}\mathcal{D}_{i},\ {\ \mathcal{D}}_{m}\cap\mathcal{D}_{n}=\emptyset,\ \forall\ m\neq n$, where $i$ is the index of a client and $\mathcal{D}_{i}$ denotes the local dataset of a client. Each client stacks the local dataset to generate the local data matrix $\textbf{X}^{(i)}$ and the local label matrix $\textbf{Y}^{(i)}$ like section II-B. Then, in the arbitrary $l$-th hidden layer, each client encodes $\textbf{Y}^{(i)}$ with a same global $\textbf{Q}_{l}$ to generate a local pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}=\textbf{Y}^{(i)}\textbf{Q}_{l}$. Specially, $\bar{\textbf{Z}}_{l}^{(i)}=\textbf{Y}^{(i)}$ when the $l$-th hidden layer is the output layer. With the real input matrix at the $l$-th layer $\textbf{X}_{l}^{(i)}$ and the pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}$, we can formulate the distributed/federated optimization problem at the arbitrary $l$-th layer as:
@@ III-A Distributed LS Problem
For convenience, we define the sum results of the two facts as $\textbf{A}_{0}$ and $\textbf{A}_{1}$ respectively in the rest of the paper. Based on the distributed LS problem, we propose the federated implementation of ACnnL, i.e. FedACnnL, in the next subsection.
III-B Federated Implementation of ACnnL
@@
Weight-Invariant Property. Based on the facts in the equation 4 and 6, we can find that the $\textbf{A}_{0}$ and $\textbf{A}_{1}$ in the equation 5 remain invariant unless the overall dataset formed by all local datasets of the clients, or the set of label encoding matrices $\mathcal{Q}$ changes. Therefore, no matter how the overall dataset is distributed among the clients, no matter what number the batch size is, and no matter how the model weights are initialized on the clients, the final closed-form LS solutions of each layer will not be changed in FedACnnL.
III-C Adaptive Batch Size Algorithm
@@
In the real-world FL scenario, the computing power and the network conditions of the clients are various. Thus, the overall time consumption of the clients completing the local computation and uploading the aforementioned resultant two items are inconsistent. This inconsistency is the straggler problem of FL, in which the PS must keep idle until the resultant items of the client with the poorest computing power and network condition arrive at the PS. An empirical practice to combat this problem is to increase the local batch size of the lagging clients to fully exert the parallelism of the hardware (CPUs or GPUs), thereby reducing the local computation time of the lagging clients. As a result, the item arrival time of the lagging clients at the PS can be advanced, reducing the idle duration of the PS. In our design, the PS adaptively adjusts the local batch size for each client in each training round by computing the ratio of each client’s time consumption to the global minimum one. The adjusted result is then sent back to each client with the trained global model weights obtained in the aggregation step. This adjustment can be mathematically represented as:
$B_{l+1}^{(i)}=min(\lceil\frac{T_{l}^{(i)}-T_{l-1}^{(i)}}{min(\{T_{l}^{(i)}-T_{l-1}^{(i)}\}_{i=1}^{C})}\times{B_{l}^{(i)}}\rceil,B_{max}^{(i)})$ (7)
@@
where $T_{l}^{(i)}$ denotes the item arrival time of the $i$-th client at the PS in the $l$-th training round and $T_{0}^{(i)}=0$. The time interval between two consecutive item arrivals of the $i$-th client, i.e. $T_{l}^{(i)}-T_{l-1}^{(i)}$ is an approximation of the client’s overall time consumption (including computation and communication) in the $l$-th training round. $B_{max}^{(i)}$ is the largest local batch size permitted to the $i$-th client and is equal to $|\mathcal{D}_{i}|$. Notably, the local batch size of each client is adjusted in every training round due to the dynamic network condition. Last but not least, the adjustment of the local batch size for each client cannot change the final model weights of each layer on each client, thanks to the weight-invariant property. Therefore, the introduction of the adaptive batch size algorithm does not influence the final model performance on each client. The pseudo-code of FedACnnL with the adaptive batch size algorithm is shown in the algorithm 1.

### v2

@@ Abstract
Analytic Federated Learning (AFL) is an enhanced gradient-free federated learning (FL) paradigm designed to accelerate training by updating the global model in a single step with closed-form least-square (LS) solutions. However, the obtained global model suffers performance degradation across clients with heterogeneous data distribution. Meta-learning is a common approach to tackle this problem by delivering personalized local models for individual clients. Yet, integrating meta-learning with AFL presents significant challenges: First, conventional AFL frameworks cannot support deep neural network (DNN) training which can influence the fast adaption capability of meta-learning for complex FL tasks. Second, the existing meta-learning method requires gradient information, which is not involved in AFL. To overcome the first challenge, we propose an AFL framework, namely FedACnnL, in which a layer-wise DNN collaborative training method is designed by modeling the training of each layer as a distributed LS problem. For the second challenge, we further propose an analytic personalized federated meta-learning framework, namely pFedACnnL. It generates a personalized model for each client by analytically solving a local objective which bridges the gap between the global model and the individual data distribution. FedACnnL is theoretically proven to require significantly shorter training time than the conventional FL frameworks on DNN training while the reduction ratio is $83\%\sim 99\%$ in the experiment. Meanwhile, pFedACnnL excels at test accuracy with the vanilla FedACnnL by $4\%\sim 8\%$ and it achieves state-of-the-art (SOTA) model performance in most cases of convex and non-convex settings compared with previous SOTA frameworks.
keywords:
[…]
1 Introduction
@@
Typical FL training schemes include first- [2, 3] (e.g. SGD) and second-order methods [4], which necessitate the computation of gradients and Hessian matrices. Despite the efficiency of these gradient descent (GD)-based methods, their reliance on gradient computations can pose challenges in practice. Specifically, (i) edge devices with poor computational capabilities are incapable of performing frequent gradient calculations. For instance, mobile phones or IoT sensors are primarily designed for inference-only purposes rather than backpropagation [5]. In such cases, only small, pre-trained or pruned models are computationally affordable for execution on these devices [1, 6]. (ii) Gradient information is not always available to support gradient computations. For instance, in situations such as federated hyperparameter tuning [7] and federated black-box attack [8], the gradients cannot be derived due to the non-differentiable nature of the loss function. As a result, GD-based FL frameworks may become infeasible in these scenarios, prompting a strong demand from both academia and industry for the development of gradient-free FL training approaches.
@@
In AFL, the global model is updated only once by using closed-form least-square (LS) solutions. Therefore, AFL can achieve a significantly shorter training time than the conventional zeroth-order FL frameworks. However, it overlooks the performance limitation of training a single global model for clients with heterogeneous data distributions. In other words, training a unified global model to satisfy the performance requirements of all clients is challenging due to the distinct preferences in data generation or collection by different clients. Personalized FL (PFL) [12] is a promising paradigm to alleviate the heterogeneous data distribution problem. In PFL, each client utilizes the common knowledge, which is the aggregated global model on the PS, to generate a personalized model tailored for its local data distribution. Meta-learning [13, 14, 15], as a representative model optimization technique in PFL, enables the aggregated global model (i.e. the meta-model) to fast adapt to heterogeneous tasks by performing several gradient descents on the global model. This motivates us to introduce meta-learning to AFL to address the heterogeneous data distribution problem in AFL. This introduction results in two challenges: First of all, the current AFL framework [11] does not have the capability of training deep neural networks (DNN). This drawback can influence the fast adaption capability of meta-learning for complex machine learning tasks. Secondly, the existing meta-learning method is only suitable for clients using GD-based optimization methods. Therefore, combining meta-learning with AFL directly is impractical since AFL does not involve any gradient approximation or computations.
@@
In this article, we first resort to ACnnL (analytic convolutional neural network learning) [16], a single-device analytic learning method for DNN, and propose its federated implementation, called FedACnnL, to enable DNN training in AFL. Specifically, we model the training of each layer as a distributed LS problem and design a layer-wise global model update rule using closed-form LS solutions to achieve a layer-wise DNN training process identical to that of ACnnL, under the federated setting. Inheriting from FedACnnL, we propose an analytic personalized federated meta-learning (APFML) framework, called pFedACnnL. A gradient-free meta-learning method is designed in it to combat the heterogeneous data distribution problem. Specifically, the PS groups the clients with similar data distribution and a meta-model that contains common knowledge is generated within each group by following the standard training process of FedACnnL. Next, each client can achieve its personalized model tailored for its local data distribution without using any gradient information by analytically solving a local PFL objective. This objective bridges the gap between the personal knowledge of its local dataset and the common knowledge of its meta-model. Our contributions can be summarized as follows:
@@
  • We first resort to ACnnL and propose FedACnnL to enable DNN training in AFL. With a competitive test accuracy of the trained DNN model, FedACnnL is proven to require a greatly shorter training time than the conventional zeroth-order FL frameworks both theoretically and experimentally.
@@
  • We validate the efficiency of our frameworks on a CPU-cluster-based testbed, which can emulate the environmental settings of data heterogeneity and resource heterogeneity in FL simultaneously. Experimentally, FedACnnL can reduce the training time by $83\%\sim 99\%$ compared with the conventional first- and zeroth-order FL frameworks. Meanwhile, pFedACnnL achieves state-of-the-art (SOTA) model performance in most cases of convex and non-convex settings, compared with other previous SOTA first- and zeroth-order PFL frameworks.
2 Preliminaries
2.1 Notations
[…]
Table 1: Table of partial notations.
@@ 2.1 Notations
2.2 Analytic Convolutional Neural Network Learning (ACnnL)
@@
Given an $N$-sample dataset $\mathcal{S}=\{\textbf{x}_{n},\textbf{y}_{n}\}_{n=1}^{N}$ where $\textbf{x}_{n}\in{\mathbb{R}}^{d_{0}}$ and $\textbf{y}_{n}\in{\mathbb{R}}^{d_{y}}$ denote the data vector and the label vector of the $i$-th sample respectively. We stack these samples to generate the data matrix $\textbf{X}=\left[\textbf{x}_{1};\textbf{x}_{2};\cdots;\textbf{x}_{N}\right]\in\mathbb{R}^{N\times{d_{0}}}$ and the label matrix $\textbf{Y}=\left[\textbf{y}_{1};\textbf{y}_{2};\cdots;\textbf{y}_{N}\right]\in\mathbb{R}^{N\times{d_{y}}}$. In the $l$-th iteration, the pseudo label matrix of the $l$-th hidden layer $\bar{\textbf{Z}}_{l}$ is generated as $\bar{\textbf{Z}}_{l}=\textbf{YQ}_{l}$, where ${\textbf{Q}_{l}}\in{\mathbb{R}^{d_{y}\times{d_{l}}}}$ is the label encoding matrix which is randomly initialized from a normal distribution. The encoding is linear so that $\bar{\textbf{Z}}_{l}$ can fit the dimensionality of $\textbf{Z}_{l}\in\mathbb{R}^{N\times{d_{l}}}$, which is the real output matrix before entering the activation function of the $l$-th hidden layer $f_{l}(\cdot)$ when X is the input of the DNN model being trained. Specially, $\bar{\textbf{Z}}_{l}=\textbf{Y}$ when the output layer of the model is being trained. Then, ACnnL formulates the supervised learning of the $l$-th layer as an least-square (LS) problem with $l_{2}$ regularization:
@@
where $\textbf{X}_{l}\in\mathbb{R}^{N\times{d_{l-1}}}$ and $\textbf{W}_{l}\in\mathbb{R}^{d_{l-1}\times{d_{l}}}$ denote the real input matrix and model weights of the $l$-th hidden layer respectively. Specially, $_{l}=\textbf{X}$ X when the input layer is being trained. $\gamma$ is the regularization coefficient which is a hyperparameter to control the search space of the LS solutions. Next, we can take the closed-from LS solutions of the weights in $l$-th hidden layer by putting the derivatives with respect to $\textbf{W}_{l}$ on the objective function in the equation 1 to 0. We present the solutions at a sample level as follows:
@@ 3 Design of FedACnnL
3.1 Distributed LS Problem
Based on the supervised learning property of ACnnL in each hidden layer, we can model the weight update problem in each layer as a distributed LS problem. Specifically, suppose we have the aforementioned $N$-sample dataset $\mathcal{S}$ in the section 2.2, and $C$ clients of edge devices. The dataset is disjointly distributed onto the $C$ clients, i.e. $\mathcal{S}=\cup_{i=1}^{C}\mathcal{D}_{i},\ {\ \mathcal{D}}_{m}\cap\mathcal{D}_{n}=\emptyset,\ \forall\ m\neq n$, where $i$ is the index of a client and $\mathcal{D}_{i}$ denotes the local dataset of a client. Each client stacks the local dataset to generate the local data matrix $\textbf{X}^{(i)}$ and the local label matrix $\textbf{Y}^{(i)}$ like section 2.2. Then, in the arbitrary $l$-th hidden layer, each client encodes $\textbf{Y}^{(i)}$ with a same global $\textbf{Q}_{l}$ to generate a local pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}=\textbf{Y}^{(i)}\textbf{Q}_{l}$. Specially, $\bar{\textbf{Z}}_{l}^{(i)}=\textbf{Y}^{(i)}$ when the $l$-th hidden layer is the output layer. With the real input matrix at the $l$-th layer $\textbf{X}_{l}^{(i)}$ and the pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}$, we can formulate the distributed/federated optimization problem at the arbitrary $l$-th layer as:
@@ 3.1 Distributed LS Problem
For convenience, we define the sum results of the two facts as $\textbf{A}_{0}$ and $\textbf{A}_{1}$ respectively in the rest of the article. Based on the distributed LS problem, we propose the federated implementation of ACnnL, i.e. FedACnnL, in the next subsection.
3.2 Federated Implementation of ACnnL
@@
Weight-Invariant Property. Based on the facts in the equation 4 and 6, we can find that the $\textbf{A}_{0}$ and $\textbf{A}_{1}$ in the equation 5 remain invariant unless the overall dataset formed by all local datasets of the clients, or the set of label encoding matrices $\mathcal{Q}$ changes. Therefore, no matter how the overall dataset is distributed among the clients, no matter what number the batch size is, and no matter how the model weights are initialized on the clients, the final closed-form LS solutions of each layer will not be changed in FedACnnL. This property is utilized in the following adaptive batch size algorithm.
3.3 Adaptive Batch Size Algorithm
@@
In the real-world FL scenario, the computing power and the network conditions of the clients are various. Thus, the overall time consumption of the clients completing the local computation and uploading the aforementioned resultant two items are inconsistent. This inconsistency is the straggler phenomenon of FL, in which the PS must keep idle until the resultant items of the client with the poorest computing power and network condition arrive at the PS. An empirical practice to alleviate this phenomenon is to increase the local batch size of the lagging clients to fully exert the parallelism of the hardware (CPUs or GPUs), thereby reducing the local computation time of the lagging clients. As a result, the item arrival time of the lagging clients at the PS can be advanced, reducing the idle duration of the PS. In our design, the PS adaptively adjusts the local batch size for each client in each training round by computing the ratio of each client’s current time consumption to the global minimum one of the first round. The adjusted result is then sent back to each client with the trained global model weights obtained in the aggregation step. This adjustment can be mathematically represented as:
$B_{l+1}^{(i)}=min(\lceil\frac{T_{l}^{(i)}-T_{l-1}^{(i)}}{min(\{T_{1}^{(i)}-T_{0}^{(i)}\}_{i=1}^{C})}\times{B_{1}^{(i)}}\rceil,B_{max}^{(i)})$ (7)
@@
where $T_{l}^{(i)}$ denotes the item arrival time of the $i$-th client at the PS in the $l$-th training round and $T_{0}^{(i)}=0$. The time interval between two consecutive item arrivals of the $i$-th client, i.e. $T_{l}^{(i)}-T_{l-1}^{(i)}$ is an approximation of the client’s overall time consumption (including computation and communication) in the $l$-th training round. $B_{max}^{(i)}$ is the largest local batch size permitted to the $i$-th client with the consideration of the client’s memory limit. Notably, the local batch size of each client is adjusted in every training round due to the dynamic network condition. Last but not least, the adjustment of the local batch size for each client cannot change the final model weights of each layer on each client, thanks to the weight-invariant property. Therefore, the introduction of the adaptive batch size algorithm does not influence the final model performance on each client. The pseudo-code of FedACnnL with the adaptive batch size algorithm is shown in the algorithm 1.